OpenStudy (adi3):

will medal Please help. find the inverse of f(n) = 2(n-2)^3

OpenStudy (adi3):

@Jhannybean

OpenStudy (adi3):

@AaronAndyson

OpenStudy (adi3):

@ganeshie8

OpenStudy (irishboy123):

\(y = 2(n-2)^3 \) \((\dfrac{y}{2})^\frac{1}{3} = n-2\) etc

OpenStudy (adi3):

how did you do that

OpenStudy (irishboy123):

\(f(n) = 2(n-2)^3\) \(y = 2(n-2)^3\) yes?

OpenStudy (irishboy123):

\(\frac{y}{2} = (n-2)^3\) ok?

OpenStudy (adi3):

why would you do y/2, isn't it